a body of mass 1kg is projected from ground at an angle of 30 with horizontal on level ground with speed 50m/s. the magnitude of change in momentum is ?
 
 

Grade:12th pass

3 Answers

Amogh Gajare
8 years ago
Change in momentum is nothing but m(v-u)
where m=mass,
u=initial velocity,
v=final velocity
This is a case of projectile motion.So,final velocity must be zero and initial velocity is given as 50 m/s
Therefore, 1(0-50)=-50 (negative because particle is decelerating)
Sourav
6 years ago
In x axis there is no change of momentum as the velocity is not changing. But in y axis the velocity is changing from "usin30°" to "-usin30°" So, change in momentum = (musin30°) - (-musin30°) = 2musin30° = 2*1*50*1/2 = 50 kg m/s
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your answer.
 
In x axis there is no change of momentum as the velocity is not changing. But in y axis the velocity is changing from "usin30°" to "-usin30°" So, change in momentum = (musin30°) - (-musin30°) = 2musin30° = 2*1*50*1/2 = 50 kg m/s
 
Thanks and regards,
Kushagra